Matheson Electronics has just developed a new electronic device that it believes will have broad market appeal. The company has performed marketing and cost studies that revealed the following information:

  1. New equipment would have to be acquired to produce the device. The equipment would cost $138,000 and have a six-year useful life. After six years, it would have a salvage value of about $24,000.
  2. Sales in units over the next six years are projected to be as follows:

Year

Sales in Units

1

7,000

2

12,000

3

14,000

4–6

16,000

  1. Production and sales of the device would require working capital of $46,000 to finance accounts receivable, inventories, and day-to-day cash needs. This working capital would be released at the end of the project’s life.
  2. The devices would sell for $55 each; variable costs for production, administration, and sales would be $35 per unit.
  3. Fixed costs for salaries, maintenance, property taxes, insurance, and straight-line depreciation on the equipment would total $149,000 per year. (Depreciation is based on cost less salvage value.)
  4. To gain rapid entry into the market, the company would have to advertise heavily. The advertising costs would be:

Year

Amount of Yearly
Advertising

1–2

$

75,000

3

$

55,000

4–6

$

45,000

  1. The company’s required rate of return is 13%.

Required:

1. Compute the net cash inflow (incremental contribution margin minus incremental fixed expenses) anticipated from sale of the device for each year over the next six years.

2-a. Using the data computed in (1) above and other data provided in the problem, determine the net present value of the proposed investment.

2-b. Would you recommend that Matheson accept the device as a new product?
Answer #1

1.

Compute the net cash inflow anticipated from sale of the device for each year as follows:

Year 1 2 3 4 5 6
a Incremental sales in units 7000 12000 14000 16000 16000 16000
b Selling price 55 55 55 55 55 55
c = a x b Incremental sales revenue 385000 660000 770000 880000 880000 880000
d = a x 35 Incremental variable costs 245000 420000 490000 560000 560000 560000
e = c - d Incremental contribution margin 140000 240000 280000 320000 320000 320000
f Incremental fixed costs 149000 149000 149000 149000 149000 149000
g Advertising costs 75000 75000 55000 45000 45000 45000
h = e - f - g Operating income -84000 16000 76000 126000 126000 126000
i Depreciation 19000 19000 19000 19000 19000 19000
j = h + i Net cash inflow from sale of device -65000 35000 95000 145000 145000 145000
Depreciation = [Cost - Salvage value]/Useful life = ($138,000 - $24,000)/6 = $19,000

2-a.

Compute the net present value of the proposed investment as follows:

Year 0 1 2 3 4 5 6
Cost of equipment -138000
Working capital employed -46000
Net cash flow from sale of device -65000 35000 95000 145000 145000 145000
Salvage value 24000
Release of working capital 46000
Net cash flows -184000 -65000 35000 95000 145000 145000 215000
Discount factor @ 13% 1.000 0.885 0.783 0.693 0.613 0.543 0.480
Present value -184000 -57525 27405 65835 88885 78735 103200
Net present value 122535

2-b.

Yes, Matheson accept the device as a new product because the net present value of the proposed investment is positive.